doggofroyo:

During a hockey game a ice surfer travel 0.75 miles what fraction represents this distance

doggofroyo:

btw,this is not on a quiz.

Hoodmemes:

U are offline but that doesn't mean that u can't come back online and see this question answered. First, as we can see we have the decimal (0.75) Wanna turn-(Convert) it into \[0.75\rightarrow \frac{ 0.75 }{ 1 }\] U wanna make it a complete fraction-(NO DECIMAL) To do this you would simply multiply each number by (100) \[0.75\times100=75\] \[1\times100=100\] Basically the broken down version of : \[\frac{ 0.75 }{ 1 }\times100=\frac{ 75 }{ 100 }\] U want to find a big number you can divide (75) and (100) by-(GCF-Greatest Common Factor) That number would be (25) \[75\div25=3\] \[100\div25=4\] The long version/broken up version of : \[\frac{ 75 }{ 100 }\div25=\frac{ 3 }{ 4 }\] So with all that being said a done the fraction that represents the distance which is (0.75 miles) is : \[\frac{ 3 }{ 4 }\] (Sorry for the long explaining,just want to make sure you fully understand the process)